A marble is selected at random from a jar containing 3 red marbles, 2 yellow marbles and 5 green marbles. What are the odds of selecting ad red marble?
{ } to { }

Answer by jessica43(140)

In probability, the odds in favor of an event or a proposition are the quantity p/(1-p), where p is the probability of the event or proposition.
So in this problem, we want p to be the probability of drawing a red marble.
This probability is the number of red marbles divided by the total number of marbles:
p=3/(3+2+5)
p=3/10
Now plug p into the above equation to find the odds:
p/(1-p)
=(3/10)/(1-(3/10))
=(3/10)/(7/10)
=3/7
So your odds are 3 to 7.